OpenStudy (anonymous):

Antiderivative of 1/x^2

OpenStudy (anonymous):

1/(x^2) = x^(-2) hence, antiderivative = x^(-1)/ -1 = -x^(-1) or -1/x + c c = integration constant

OpenStudy (lgbasallote):

change it to negative exponent then put it on top then integrate by power rule

OpenStudy (anonymous):

THANK YOU:D

OpenStudy (anonymous):

What about 1/2x+3? Im confused about this one

OpenStudy (lgbasallote):

this is all yours @Rohangrr

OpenStudy (anonymous):

@lgbasallote wat??

OpenStudy (inkyvoyd):

He's saying, don't you copy paste ever again.

OpenStudy (lgbasallote):

btw..while you're at it there's this problem confusing me...a quite similar one...how to integrate \(\LARGE \int e^{x^{-2}} dx\) i think it has something to do with u-substitution as well....

OpenStudy (anonymous):

@inkyvoyd just say quiet if you aren't answering

OpenStudy (mimi_x3):

igbiw that's an error function..

OpenStudy (inkyvoyd):

I encourage you guys to report this blatant plagiarizer http://answers.yahoo.com/question/index?qid=20090119012032AAtrnFc.

OpenStudy (lgbasallote):

oh....no wonder i cant get it :/

OpenStudy (inkyvoyd):

@lgbasallote , stop faking xD

OpenStudy (mimi_x3):

i think that is thhe one used in statistics.

OpenStudy (lgbasallote):

hahahaha =))))) well i really couldnt get it :P and btw...the question was deleted...

OpenStudy (inkyvoyd):

@Mimi_x3 , it's used in a lot of things

OpenStudy (kinggeorge):

This link works though http://answers.yahoo.com/question/index?qid=20090119012032AAtrnFc

OpenStudy (lgbasallote):

"1/(x^2) = x^(-2) hence, antiderivative = x^(-1)/ -1 = -x^(-1) or -1/x + c c = integration constant" copied straight from it...does look familiar....

OpenStudy (lgbasallote):

(spacings included)

OpenStudy (inkyvoyd):

Huh. The question is the exact same too.

OpenStudy (lgbasallote):

no...the question had parenthesis...or did it...?